OpenStudy (anonymous):

Factor the following polynomial: x2 + 10x + 21 A. (x + 2)(x + 5) B. (x + 1)(x + 10) C. (x + 7)(x + 3) D. (x + 21)(x + 1)

OpenStudy (anonymous):

Check your answer choices by simplifying them. Also, factoring is finding \(a\) and \(b\) in \((x+a)(x+b)=x^2+(a+b)x+ab\) such that \(a+b=10\) and \(ab=21.\)

OpenStudy (anonymous):

Find numbers that have a product of 21 and a sum of 10. :D

mathslover (mathslover):

@anhhuyalex You are requested to not post such type of comments .. come to the point and then explain the asker ... dont confuse him/her more and more

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

Pardon me?

mathslover (mathslover):

No problem @anhhuyalex this is great to see that u accepted ur mistake

OpenStudy (anonymous):

\[x ^{2} +3x + 7x + 21 =(x+3)(x+7) \] So option C is correct.If You meant what I mean.

OpenStudy (anonymous):

@anhhuyalex, maths is trying to say that you should not repeat what has been said as it causes confusion.

OpenStudy (anonymous):

I'm really confused. :))

OpenStudy (anonymous):

Oh. I posted about 2 seconds after you did. But of course I apologize :D

OpenStudy (anonymous):

is the answer C ?

OpenStudy (anonymous):

@anhhuyalex, it is perfectly fine. The idea is that, when someone is just learning something, it's not obvious that the same thing is being said in different words. So, when you have only one person explaining, it prevents confusion. @devarsh.ruparelia, you are not supposed to post the answers here. Read the Code of Conduct. Thanks.

OpenStudy (anonymous):

@Limitless Thanks for your advise .I didnt know of this .From now on I will just give the person a help or method to solve problem.

OpenStudy (anonymous):

@devarsh.ruparelia, no. Thank you, sir! It is not often that people respect the rules here. I am very happy you have chosen to. :)

OpenStudy (anonymous):

(Sorry if you're female. I can't read gender in your name properly.)

OpenStudy (anonymous):

@Limitless I am male.